If y = sin(ex), then y’ is

Select one:
A. excos(ex)
B. xcosx
C. cos(ex)
D. xcosx-1
1
Expert's answer
2020-10-22T17:59:21-0400

Given ,

"y=sin(e^x)"

Differentiate y With respect to x using chain rule

"\to y'=cos(e^x)\times\dfrac{d(e^x)}{dx}"


"\to y'=cos(e^x)\times e^x"


"\to y'=e^xcos(e^x)"


So option (A) is correct.